OpenStudy (anonymous):

How would i Simplify 7+i / 2-2i (Complex Numbers)

OpenStudy (ivettef365):

You have 7+i 2+2i ----- . ------- 2-2i 2+2i

OpenStudy (anonymous):

Then i would now just combine like terms and foil correct?

OpenStudy (ivettef365):

right and remember that i^2 = -1

OpenStudy (anonymous):

ok Thakns!

OpenStudy (ivettef365):

Like when you multiply (2-2i)*(2+2i) =4-4i^2 = 4-4(-1) = 4+4 = 8

OpenStudy (anonymous):

ok after i have the step tht you just shown here, do i just keep simplifying it down?

OpenStudy (anonymous):

Speaking with this in mind: =4-4i^2=4-4(-1)=4+4=8 what is the next step

OpenStudy (ivettef365):

the step I did is the bottom part the top part as follows: (7+i) *(2+2i) 14+14i+2i+2i^2 14 +16i +2(-1) = 16i + 14-2 16i+12

OpenStudy (ivettef365):

Now you have 16i + 12 --- ---- 8 8

OpenStudy (ivettef365):

Now just simplify

OpenStudy (anonymous):

16i + 12 ---- ------ 8 8 that wuld be divide each by 8 correct?

OpenStudy (anonymous):

if so , after doing tht i got 8i + 1.5 does not seem correct

OpenStudy (ivettef365):

16/8 = 2, so it would be 2i+3/2

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (ivettef365):

your welcome :)